UT 20140121 LUCI2 Nighttime Commissioning

Observer: JHeidt, APramskiy, DThompson
Telescope Operator: GBechetti
LBTO Support: DMiller (for AO-preset)

Plan for the night

- AO-preset
- Pointing model
- Deep imaging MOS/Telluric tests
- Ima movement due to flexure

Summary

Daten taken for a new pointing model

LS/MOS-spectroscopy

Preparations

Field alignment: dx,dx = 700,1400 Link communication from the DX WUnit to the DX AdSec was attempted to fix
but couldn't be established, ie AO-preset on SX only

Details

All times are UT, files are 20140120.iiii

23:03 (Jan20) Readnoise: LIR; 2.7399sec, #1-20, MER, 6.8915sec, #21-40

00:46 Sunrise, clear, some cirrus around, low wind, hum=5%, T=3deg

01:00 Geno checks pointing and collimation

01:10 AO-preset, right side, not successfull,

01:25 BS9107, several changes in RTD, check if LS-acq. still works, #46-49

01:50 Test if acq. on two frames with different RA works, throughslit SCO_ARCS_EVA @ Bs9107 #50,
point to BS9108, src #51, works!!!

02:15 HH91, FeII #61-81, seeing started with 0.6" on images, but went to 2-3" on guider and fluctuates strongly after 10mins
stopped at 03:20, images even not useful for nice images

03:45 Start with pointing model LUCi2, put in pointing mask, #82-172, after 60 stars (06:00) seeing is still poor (1.5-2")
we continue to catch all 88 stars available
### UT 20140121 
luci2.20140120.NNNN.fits
All data are 10x2.75s LIR/int

Pointing model data collection 

Br_gam, Pointing mask, N375 imaging, PA=0 TRACK mode
0082      - test exposure
0083 0084 - blind filter in FW2

03:58 @ start
Pointing log: /home/telescope/PointingLogs/20140121_*X.log
No collimating done during the pointing model data taking

Seq NNNN StarID
 0 0085 WT10_357 w/blind filter still in
 1 0086 WT10_357 Br_gam (and all subsequent)
 2 0087 WT10_409 
 3 0088 ACT0400
 4 0089 ACT0342
 5 0090 ACT0293
 6 0091 WT10_191  (unwrap AZ before this image)
 7 0092 WT10_265
 8 0093 WT10_108
 9 0094 ACT0197 
11 0095 ACT0237 
12 0096 WT10_320
13 0097 WT10_333
14 0098 WT10_288
15 0099 WT10_271
16 0100 WT10_316
10 0101 WT10_269
17 0102 WT10_214
18 0103 ACT0110 
19 0104 ACT0159 
20 0105 WT10_299
21 0106 ACT0263
22 0107 WT10_261
23 0108 ACT0295
25 0109 WT10_402
26 0110 WT10_410
27 0111 WT10_407
29 0112 WT10_372
30 0113 WT10_392
31 0114 WT10_390
28 0115 ACT0331
32 0116 ACT0410
33 0117 WT10_368
34 0118 ACT0384
35 0119 WT10_353
36 0120 ACT0355
37 0121 ACT0329
24 0122 WT10_344
39 0123 ACT0246
40 0124 WT10_174
38 0125 ACT0299 
41 0126 WT10_176
42 0127 ACT0195
43 0128 ACT0250
44 0129 WT10_197
45 0130 ACT0180
46 0131 WT10_146
47 0132 WT10_218
48 0133 WT10_271
49 0134 ACT0303
50 0135 ACT0320
51 0136 WT10_307
52 0137 WT10_350
53 0138 ACT0337
54 0139 ACT0361
55 0140 WT10_348
56 0141 WT10_318
57 0142 WT10_333
58 0143 ACT0288
59 0144 WT10_237
60 0145 WT10_184
62 0146 ACT0235
63 0147 WT10_252
64 0148 ACT0286
65 0149 WT10_269
66 0150 ACT0180
67 0151 ACT0197
68 0152 ACT0178
69 0153 ACT0142
71 0154 ACT0231
72 0155 WT10_267
73 0156 ACT0301
74 0157 ACT0282
61 0158 WT10_254
70 0159 ACT0214
75 0160 ACT0265
76 0161 WT10_314
77 0162 WT10_329
78 0163 WT10_342
79 0164 ACT0340
80 0165 WT10_377
82 0166 WT10_346
83 0167 ACT0370
81 0168 WT10_379
84 0169 ACT0388
85 0170 ACT0390
86 0171 ACT0381
88 0172 WT10_374
87 0173 - (no star available, skipped) 
Stop pointing logging

Correct pointing 
Collimate on-axis BS9134 @ mid EL
Initial collimation after 3h was not bad, star round
stop guiding and WFSing
move AGw stages out of beam (home them)
note: need to correct for internal flexure using the dot circle!

0173  PA=0    975.59 1050.58
0174  PA=180  983.19 1061.71
0175  PA=270  982.01 1052.00
0176  PA=90   970.26 1057.02
0177  image w/blind filter in

07:05 CofR using DMillers approach, #173-177

07:28 Ima movement, M67, PA=0,90,180,280, #178-181

07:44 HS1024, yak, moon to close, can't collimate

08:06 HS0911, no GS, check pointing and sent preset again, OK, sky #182, src #183, slitview #184, verification #185, spectra #186-205, first GEIRS error since 5n after #195, doesn't impact the execution of the obs

09:26 (approx) Telluric for quasar 0915+47. #207 sky (script should not have taken a sky frame), #208 src, #209 slit, #210 post-offset, #211-212 spectra of telluric star

09:46 (approx) M67. #213 @ PA=0, #214 @ PA=90, #215 @ PA=180, #216 @ PA=270

10:06 (approx) SDSS1110. #217 sky, #218 src, #219 slit: offset: +4.71" -2.74", #220 post-offset. #221-228 spectra 1x300s G150@2.245um

11:20 Another set of SDSS1110, sky #229, src #230, MOSview #231, verification #232-235, specta #236-240,
GEIRS error, #241 probably, #242 certainly crap

12:34 Telluric for SDSS1110, Geno needs to correct pointing (seems to be far way off), another GEIRS error, Alexander restarts driver
sky #243, src #244, MOSview #245/246, align, verification pos1 #247/248, spectrum #249, verification pos2 #250, spectrum #250 (3sec), #251 (4*3sec int), verification pos3 #252-256, spectrum pos3 #257, spectrum pos4 #258
yak, spectra taken with N3.75 camera

13:27 RG1647, guide star too weak...towards east

13:38 BS9142, almost photometric now, ZPs, zJHKKs #260-269

13:47 Move to zenith for flats

13:49 J_low, 10sec, 4k-27k, #270-289
13:57 Ks, 4sec, 5k #290-299
14:00 Ks, 7sec, 15K, #300-309

14:05 Closing

Calibrations

Kurk, Mask SDSSJ1110, G150@2.245mue, clear+K, CalUnit in, all images LIR, all normal
Lamp off, 5*2.75s, #310-314
Halo2 on, 5*10sec, #315-319
Lamps off, 5*10sec, #320-324
Argon on, 1*10sec, #325
Neon on, 1*10sec, #326
Xenon on, 1*10sec, #327
Lamps off, 5*10sec, #328-332
Spectro sieve, 10*10sec, G150@2.245mue, clear+K, #333-342
Spectro sieve, 10*10sec, G210@2.225mue, clear+K, #343-352

Worseck, 0.75" slit, G150@2.17mue, clear+K, CalUnit in, all images LIR, all normal
Lamp off, 5*2.75s, #353-357
Halo2 on, 5*15sec, #358-362
Lamps off, 5*15sec, #363-367
Argon on, 1*10sec, #368
Neon on, 1*10sec, #369
Xenon on, 1*10sec, #370
Lamps off, 5*10sec, #371-375

-- JochenHeidt - 21 Jan 2014
Topic revision: r8 - 21 Jan 2014, JochenHeidt
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback